A Complete Bibliography of Computing Systems Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 February 2025 Version 1.52

Title word cross-reference

4.3BSD [BL90]. 4.4BSD [McK95].

9 [PPD^{+}95, Wel94].

Access [CLS92]. Accesses [BD96]. Achieving [FC96]. across [ACO92]. Adaptation [NPS95]. Adaptive [DK95, MP90]. Address [KC96]. Adrift [DGT96]. Against [Car91]. Ahead [Win94]. Algorithm [BK91, Win94]. Algorithms [BS93, CS90]. Amoeba [DOKT91]. among [CBE93]. Animation [BK91]. Announcements [Sal94]. Applicability [KCCT94]. Application [Bis88, NPS95]. Application-Aware [NPS95]. Applications [BDWW89, PK93]. Approach [BL90, LaP94, Mor88]. approaches [SEKN92]. Archaeology [Gra92]. Architectural [Ros92].

Architecture [BBD^{+}91, BSKH93, Gol92, VLD^{+}92]. Architectures [Wel94]. Area [IP94]. Arjuna [PSWL95]. ASCII [HF91]. Assessment [SGH^{+}89]. Assignment [Win94]. Atomic [MKC92]. Attached [Puc91]. Audio [Haw90]. Authentication [KCCT94]. Autonomous [DLO92]. Availability [VF95]. Aware [NPS95].

Based [CK96, Neu92, YKB94]. Basic [CS90]. Batch [CS94]. Be [Pik88]. Bell [PPD^{+}95]. Bottlenecks [ACO92]. BSD [Ste89]. Building [Cal96b, Cal96a]. bus [OW94].

C [Sak92a, Car91, Joh88, Nil94, Par95, Sak92b, Str88, Str89c, Str89a, Str89b, Wal91]. Cached [VLD^{+}92]. caches [OW94]. Caching [KAD96, Wel91]. Calls [PA90]. Can [Les88a]. Cards [KCCT94]. Case [Car91, Wal91]. Casper [VLD^{+}92]. Checking [BD96, HT94]. Choices [CIM92]. CHORUS [RAA^{+}88]. CIA [Gra92]. Clocks [DGT96, Yah94]. Clouds [DCM^{+}90, DAM^{+}91]. code [Les88a]. Collaborative [PK93]. Collection [Nil94]. Communication [AOW96, BZM91, CK96]. Comparison [CS90, DOKT91, Wel94, SEKN92]. Compliant [KAD96]. Component [WCN93]. Components [CBE93]. Compression [Haw90]. Computer [LaP94, Wil88a, Wil88b]. Computers [DK95]. Computing [NPS95, Win96]. Concurrent [Pik89, STV90]. Conditions [BD96]. Configurable [Puc91]. Configuration [Bur95]. Configurations [Cal96b, Cal96a]. Connected [HH95]. Conservative [BL90]. Considerations [War94]. Consistency [Gol92]. Constraints [HT94]. Contributors [Ano95, Ano96a, Ano96b, Ano96c]. Control [CLS92]. Controlling [Lib91]. Controversy [Ada96, Car91, Pik88, Wal91]. Conventional [DV90]. Cooperative [CBE93]. copy [SM88]. copy-on-write [SM88]. CORBA [KAD96]. Correspondences [WCN93]. Corrigendum [Cal96a, Mof92a, Sak92a]. Cost [Yah94]. CPU [CS90]. Critique [Sak92b, Sak92a].

Data [ACO92, Bis88, CL91, Gri89, Haw90, HF96, Puc91, Sri93, War94, HF91]. database [HF91]. Dealing [WCN93]. dear [O’D94b]. Delayline [IP94]. Demultiplexing [MD92]. Design [Baa88, DCM^{+}90, Gra92, KP93, PSWL95, PHS96, Smi91, Sos95, OW94]. Developing [BDWW89, BCR91]. Development [BSKH93]. Directing [Sos95]. discovery [SEKN92]. Disk [CL91, DK95, Ste89]. Distributed [BBD^{+}91, BZM91, CL91, Car95, DLO92, DCM^{+}90, DAM^{+}91, DOKT91, Gol92, LM94, PK93, RAA^{+}88, STV90, Wag90, Wel94, Win94, WRW96]. Distribution [SSZ^{+}93, YKB94]. Documenting [CI93]. Domain [BSSW96]. Down [DK95]. drive [Ste89]. DUNE [PA90]. DUNE iX [BDKM93]. Durra [BDWW89]. Dynamics [Wil88a, Wil88b]. Dynascope [Sos95].

Economical [Mof92b]. Editing [KTBL93]. Editor [Ree95, Sal95a, Sal95b, Tho90]. Editorial [Avo96, Deu92, Dew93, Hon94, Pre95, Sch96, Spa89, Spa91, UB93]. Editors [KP93]. Effects [SM88]. Efficient [MD92]. Egg [Joh93]. Electronic [PHS96]. Elk [LB94]. Emulation [IP94]. Encryption [Bis88]. Enforcement [BSSW96]. Engine [Bur95]. Engineering [MBM93, Mor88]. Enhanced [BW88]. Environment [CBE93, KTBL93, Wag90]. Environments [BSKH93, Sri93]. Erratum [Wil88b]. Evaluation [BS93, SSZ^{+}93]. Evolution [BZM91, Str89a]. Exeunt [Sal96]. expect [Lib91]. Expected [KC96]. Experience [BCR91, Duf89]. Experiences [ACO92]. Experimental [CDM90]. Extending [BP88]. Extensible [Tho90]. Extension [LB94].

Facility [MKC92]. Fall [Joh93]. Fast

[Bis88, HF91]. Features [DJ88]. Feedback [MP90]. Felicitations [O’D88d]. File [BP88, BD96, CS94, FC96, Neu92, Wel91, Wel94]. Files [Mof92b]. Filesystem [McK95]. Fine [CLS92, MP90]. Fine-Grain [MP90]. Fine-Grained [CLS92]. Firewall [Cal96b, Cal96a]. Firewall-Router [Cal96b, Cal96a]. fork [SM88]. Formal [LaP94]. Formatter [KV89]. Framework [CI93, PHS96]. Frameworks [CIM92]. friends [O’D94b].

Garbage [Nil94]. Generation [Par95]. Global [Neu92]. Golden [Joh93]. Goose [Joh93]. GRAB [Les88b]. Grain [MP90]. Grained [CLS92]. Graphics [Tho90, Wil88a, Wil88b]. Grasshopper [DdBF^{+}94]. Greetings [Bis94, O’D88b, O’D88d, O’D88c, O’D89a, O’D89b, O’D89c, O’D89d, OS89, O’D90a, O’D90b, O’D90c, O’D91a, O’D91b, OS91, O’D91c, O’D92b, O’D92c, O’D92d, O’D92a, O’D93a, O’D93b, OS94, O’D94a, O’D94b, O’D88a]. Group [KP93]. GroupDesign [KTBL93]. Guest [Avo96, Deu92, Dew93, Hon94, Ree95, Sch96, Spa89, Spa91, UB93]. Guide [BBD^{+}91].

Heterogeneous [BDWW89, CBE93, KTBL93, Sri93, Wag90]. Heuristics [Ste89]. High [CL91, CK96, PHS96]. High-Performance [CK96]. High-Speed [PHS96]. Hints [PA90]. History [DGT96]. Hypertext [Bro89].

I/O [CL91]. Icon [Gri89]. IK [SSZ^{+}93]. Imaging [PHS96]. Implementation [BBD^{+}91, Bis88, CBE93, CDM90, DCM^{+}90, DGT96, PSWL95, SLM^{+}90, Sos95, Baa88]. Implementing [MKC92]. Incoming [MD92]. Independent [Ada96, Rub96a, Rub96b]. Index [Ano93, Ano96d]. Indexes [Les88b]. Indexing [DLO92]. Information [Gol92, VF95]. Inheritance [Car91, Sak92b, Sak92a, Wal91, Str89b]. Instruction [Mor88]. Integration [WCN93]. Intelligent [FC96]. Inter [WCN93]. Inter-relation [WCN93]. Interactive [Lib91]. Interdependent [Win94]. Interface [McK95, NPS95]. Interfaces [ACO92]. Internet [DLO92, SEKN92]. Interoperable [Sri93]. Interrupt [Col96]. Intuitive [HF96]. Inversion [Mof92b]. Inverted [Les88b]. Issue [Ano95, Ano96a, Ano96b, Ano96c]. Issues [KP93, SLM^{+}90].

Java [RWWB96, WRW96].

Kerberos [DGT96]. Kernel

[PMI88, Rua90]. Keynote [Tho90]. Killing [Joh93]. Kit [LB94].

Laboratory [Smi91]. Labs [PPD^{+}95]. Laid [Joh93]. Language [Car95, DJ88, Gri89, LB94, Tho90]. Languages [Lan90]. Large [Mof92b]. last [O’D94b]. Latency [VF95]. level [OW94]. Leveraging [CK96]. Lifetime [KC96]. Linkage [Str88]. Little [Lan90]. Log [CS94]. Log-Structured [CS94]. Longer [FG90]. Look [Win94]. Look-Ahead [Win94]. Low [Les88b].

Mach [BL90]. Mach/4.3BSD [BL90]. Machine [BDWW89]. make [Baa88]. Makeup [KV89]. Management [BS93, SM88]. Managing [Sal95a, Sal95b, Sri93]. Manipulation [Puc91]. Measured [Wel91]. Mechanism [AOW96]. Mechanisms [CBE93]. Medical [PHS96]. meets [Joh88]. megabyte [OW94]. Memory [CK96, SM88, OW94]. Memory-Based [CK96]. Messages [AOW96]. Messaging [CK96]. Migration [CCK^{+}95]. minimizing [OW94]. Mobile [DK95, NPS95]. Model [DV90, Neu92, WRW96]. Modeling [LaP94]. MPVM [CCK^{+}95]. Multi [BSKH93]. Multi-User [BSKH93]. Multidatabase [BS93, CBE93]. Multiple [Car91, KP93, Str89b, Wal91]. Multiprocessor [CS90, Puc91, SLM^{+}90]. multiprocessors [OW94]. Music [Lan90, Tho90].

Naming [CDM90]. Navigation [YKB94]. Network [BDWW89, IP94, KCCT94, Wel91]. Networks [BDWW89, War94]. NFS [Mog94]. No [FG90]. Non [Win96, Yah94].

Non-Synchronized [Yah94]. Non-Technical [Win96]. Note [Sal95a, Sal95b].

O [CL91]. Object [AOW96, BBD^{+}91, CLS92, CI93, DV90, Gra92, KAD96, PHS96, SGH^{+}89, WRW96]. Object-Oriented [AOW96, BBD^{+}91, CLS92, CI93, Gra92, PHS96, SGH^{+}89]. Objects [DAM^{+}91, MKC92]. Omnes [Sal96]. One [Ada96, Rub96a, Rub96b, O’D94b]. One-Time [Ada96, Rub96a, Rub96b]. Operating [BDKM93, BCR91, DCM^{+}90, DdBF^{+}94, DV90, DJ88, KC96, Ros92, RAA^{+}88, SLM^{+}90, SGH^{+}89]. Operation [HH95]. Operations [SM88]. Optimal [CS94, OW94]. Optimized [CK96]. Orchestra [Haw90]. Oriented [AOW96, BBD^{+}91, CLS92, CI93, Gra92, PHS96, SGH^{+}89]. Orthogonal [Ros92]. Orthogonally [DdBF^{+}94]. Output [KV89]. Overcoming [ACO92]. Overhead [Les88b]. Overview [SSZ^{+}93].

Packets [MD92]. Page [KV89]. Parallel [AOW96, Baa88]. Parallelization [BL90]. Parametrized [Str89c]. Partially [HH95]. Passwords [Ada96, Rub96a, Rub96b]. Performance [BS93, CK96, PHS96, Wel91]. Persistence [Ros92, SSZ^{+}93, VLD^{+}92]. Persistent [DdBF^{+}94]. Personal [Haw90]. Perspectives [SGH^{+}89]. Pickling [RWWB96]. Plan [PPD^{+}95, Wel94]. Platform [SSZ^{+}93]. Policies [DK95]. Policy [War94]. Portability [FG90]. positioning [Ste89]. Postprocessing [KV89]. Predictable [FC96]. Principles [Sak92b, Sak92a]. Priorities [Col96]. Problem [FG90]. Procedure [PA90]. Process [Rua90]. Processes [Lib91]. Processing [BZM91]. Programming [DAM^{+}91, DJ88, Gri89, NPS95, STV90]. Prospero [Neu92]. Protocols [DGT96].

Prototype [BSSW96]. Provide [CL91]. Psyche [SLM^{+}90]. PVM [CCK^{+}95].

queries [HF91]. Querying [WCN93].

Queueing [Col96].

Race [BD96]. Radix [MBM93]. RAID [LM94, BZM91]. Rates [CL91]. Real [BDKM93, DJ88, Nil94]. Real-Time [BDKM93, Nil94, DJ88]. Recovery [Mog94]. Redundancy [Sri93]. Refinement [CIM92]. Rejoinder [Ada96]. relation [WCN93]. Relational [HF96, WCN93]. Relations [WCN93]. relatively [HF91]. RelaX [MKC92]. Reliable [Nil94]. Remote [PA90]. Representing [HF96]. Resource [BW88, SEKN92]. Response [FC96, Rub96b, SM88]. Router [Cal96b, Cal96a]. RP3 [BCR91]. Rule [LaP94]. Rule-Set [LaP94].

safe [Str88]. Scheduling [CS90, MP90]. Schema [HF96, WCN93]. Scope [Car95]. Scripts [Lib91]. second [OW94]. second-level [OW94]. secret [Les88a]. Secure [Yah94]. Security [HT94]. Server [Sos95]. Servers [VF95]. Services [DLO92, Gol92]. Set [LaP94]. Setting [Col96]. Shared [KTBL93, OW94]. shared-bus [OW94]. shared-memory [OW94]. Sharing [BW88]. Should [Pik88]. Single [KC96]. Single-Address-Space [KC96]. Site [Bur95]. Size [CS94]. Smart [AOW96, KCCT94]. Software [BSKH93, Col96, Mor88, Smi91]. Solved [FG90]. Sort [MBM93]. SOS [SGH^{+}89]. source [Les88a]. Space [KC96]. Special [Bis94, OS94]. Specifying [HT94]. Speed [PHS96]. Spin [DK95]. Spin-Down [DK95]. Spooling [Wag90]. Sprite [DOKT91, Wel91, Wel94]. Spritely [Mog94]. stable [HF91]. Standard [Bis88]. State [RWWB96]. Storage [Les88b]. Striping [CL91]. Strong [Sal90]. Structure [HF96].

Structured [CS94]. Structures [Gri89]. Stub [Par95]. SUN [ACO92]. SUN-Transputer [ACO92]. Support [Ros92, STV90]. Supporting [KP93, VLD^{+}92]. survive [Les88a]. Swift [CL91, LM94]. Swift/RAID [LM94]. Synchronization [Rua90]. Synchronized [Yah94]. Synchronous [PK93]. Synthesis [PMI88]. System [BBD^{+}91, BDKM93, BK91, BP88, BZM91, Bro89, BCR91, CLS92, CI93, CBE93, CK96, CDM90, DCM^{+}90, DAM^{+}91, DdBF^{+}94, DJ88, FC96, KAD96, LaP94, LM94, Neu92, Par95, Pik89, RWWB96, Ros92, RAA^{+}88, SLM^{+}90, SGH^{+}89, Wel91, Wel94, WRW96, McI89]. Systems [AOW96, CS94, DV90, DOKT91, Duf89, KC96, Pik88, STV90, Win96, YKB94].

Tasks [Win94]. TCP [MD92]. Technical [Win96]. Technique [CI93]. Text [KV89, Mof92b]. Threads [DAM^{+}91]. Threat [Win96]. Three [Wel94]. Tilde [CDM90]. Time [Ada96, BDKM93, FC96, Nil94, Rub96a, Rub96b, SM88, DJ88, O’D94b]. Timeliness [Yah94]. Tom [Sal90]. Tool [Cal96b, Cal96a, IP94]. Toolkit [KP93, PK93]. traffic [OW94]. Transaction [BZM91, BS93, MKC92]. Transactional [CLS92]. Transfer [ACO92]. Transparent [CCK^{+}95, Pik88]. Transputer [ACO92]. Trees [HF96]. Trust [YKB94]. Trust-Based [YKB94]. Trusted [LaP94]. Tuple [HF96]. Tutorial [Wil88a]. Two [BS93, DOKT91]. Type [BSSW96, Str88]. Type-safe [Str88]. Types [Str89c].

Undergraduate [Mor88]. United [Joh93]. UNIX [BSSW96, BW88, BP88, Bro89, CS90, Duf89, HT94, McI89, PK93, SM88, Les88a]. Unorthodox [Mor88]. Updater [FC96]. User [BSKH93, KCCT94]. Using [CL91, PA90, MP90]. UTS [Rua90].

Version [CCK^{+}95]. via [Col96]. Virology [McI89]. Virtual [McK95, Neu92]. virus [McI89]. Viruses [Duf89]. Vnode [Wel94]. Volumes [Ano96d, Ano93].

Watchdogs [BP88]. Weak [Gol92].

Weak-Consistency [Gol92]. Web [VF95]. Welcome [Nem88]. Wide [IP94, VF95]. Wide-Area [IP94]. Window [Pik88, Pik89]. Workload [Win94]. World [VF95]. Write [CS94, SM88].

YACC [Joh88].

References

Atkins:COMPSYS-5-2-159

[ACO92] M. Stella Atkins, Yan Chen, and Florina Olariu. Experiences: Overcoming data transfer bottlenecks across SUN-Transputer interfaces. Computing Systems, 5 (2):159–192, Spring 1992. CO-DEN CMSYE2. ISSN 0895-6340.

Adams:COMPSYS-9-3-255

[Ada96] John Adams. Controversy: Rejoinder: Independent one-time passwords. Computing Systems, 9(3):255–256, Summer 1996. CO-DEN CMSYE2. ISSN 0895-6340. See [Rub96a, Rub96b].

Anonymous:COMPSYS-6-1-37

[Ano93] Anonymous. Index volumes 1– 5 (1988–1992). Computing Systems, 6(1):37–50, Winter 1993. CODEN CMSYE2. ISSN 0895-

Anonymous:COMPSYS-8-4-415

[Ano95] Anonymous. Contributors to this issue. Computing Systems, 8(4): 415–416, Fall 1995. CODEN CM-SYE2. ISSN 0895-6340.

Anonymous:COMPSYS-9-1-85

[Ano96a] Anonymous. Contributors to this issue. Computing Systems, 9(1): 85–87, Winter 1996. CODEN CMSYE2. ISSN 0895-6340.

Anonymous:COMPSYS-9-3-259

[Ano96b] Anonymous. Contributors to this issue. Computing Systems, 9(3): 259–260, Summer 1996. CODEN CMSYE2. ISSN 0895-6340.

Anonymous:COMPSYS-9-4-407

[Ano96c] Anonymous. Contributors to this issue. Computing Systems, 9(4): 407–409, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Anonymous:COMPSYS-9-4-413

[Ano96d] Anonymous. Index to volumes 1– 9 (1988–1996). Computing Systems, 9(4):413–428, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Arjomandi:COMPSYS-9-4-313

[AOW96] Eshrat Arjomandi, William G.

O’Farrell, and Gregory V. Wilson. Smart messages: An objectoriented communication mechanism for parallel systems. Computing Systems, 9(4):313–329, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Avolio:COMPSYS-9-1-1

[Avo96] Frederick M. Avolio. Guest editorial. Computing Systems, 9(1):

1–2, Winter 1996. CODEN CM-SYE2. ISSN 0895-6340.

Baalbergen:COMPSYS-1-2-135

[Baa88] E. H. Baalbergen. Design and implementation of parallel make. Computing Systems, 1(2):135– 158, Spring 1988. CODEN CM-SYE2. ISSN 0895-6340.

Balter:COMPSYS-4-1-31

[BBD^{+}91] R. Balter, J. Bernadat, D. Decouchant, A. Duda, A. Freyssinet, S. Krakowiak, M. Meysembourg, P. Le Dot, H. Nguyen Van, E. Paire, M. Riveill, C. Roison, X. Rousset de Pina, R. Scioville, and G. Vandôme. Architecture and implementation of guide, an object-oriented distributed system. Computing Systems, 4(1): 31–67, Winter 1991. CODEN CMSYE2. ISSN 0895-6340.

Bryant:COMPSYS-4-3-183

[BCR91] Ray Bryant, Hung-Yang Chang, and Bryan Rosenburg. Experience developing the RP3 operating system. Computing Systems, 4(3):183–216, Summer 1991. CO-DEN CMSYE2. ISSN 0895-6340.

Bishop:COMPSYS-9-2-131

[BD96] Matt Bishop and Michael Dilger. Checking for race conditions in file accesses. Computing Systems, 9(2):131–152, Spring 1996. CO-DEN CMSYE2. ISSN 0895-6340.

Banino:COMPSYS-6-4-425

[BDKM93] Jean-Serge Banino, Jean Delcoigne, Claude Kaiser, and Gérard Morisset. The DUNE iX real-time operating system. Computing Systems, 6(4):425–480, Fall 1993. CODEN CMSYE2. ISSN 0895-6340.

Barbacci:COMPSYS-2-1-7

[BDWW89] Mario B. Barbacci, Dennis L.

Doubleday, Charles B. Weinstock, and Jeannette M. Wing. Developing applications for heterogeneous machine networks: The Durra network. Computing Systems, 2(1):7–35, Winter 1989. CODEN CMSYE2. ISSN 0895-6340.

Bishop:COMPSYS-1-3-221

[Bis88] Matt Bishop. An application of a fast data encryption standard implementation. Computing Systems, 1(3):221–254, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Bishop:COMPSYS-7-1-v

[Bis94] Matthew A. Bishop. Special greetings. Computing Systems, 7 (1):v–vi, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Bentley:COMPSYS-4-1-5

[BK91] Jon L. Bentley and Brian W.

Kernighan. A system for algorithm animation. Computing Systems, 4(1):5–30, Winter 1991. CODEN CMSYE2. ISSN 0895-

6340. URL http://www.usenix.
org/publications/compsystems/
1991/win_bentley.pdf;   https:
//www.usenix.org/publications/
compsystems/computing-systems-

Boykin:COMPSYS-3-1-69

[BL90] Joseph Boykin and Alan Langerman. Mach/4.3BSD: A conservative approach to parallelization. Computing Systems, 3(1):69–99, Winter 1990. CODEN CMSYE2. ISSN 0895-6340. Bershad:COMPSYS-1-2-169

[BP88] Brian N. Bershad and C. Brian Pinkerton. Watchdogs—extending the UNIX file system. Computing Systems, 1(2):169–188, Spring 1988. CODEN CMSYE2. ISSN 0895-6340.

Brown:COMPSYS-2-1-37

[Bro89] P. J. Brown. A hypertext system for UNIX. Computing Systems, 2 (1):37–53, Winter 1989. CODEN CMSYE2. ISSN 0895-6340. Breitbart:COMPSYS-6-3-245

[BS93] Y. Breitbart and A. Silberschatz. Performance evaluation of two multidatabase transaction management algorithms. Computing Systems, 6(3):245–283, Summer 1993. CODEN CMSYE2. ISSN 0895-6340. Ben-Shaul:COMPSYS-6-2-65

[BSKH93] Israel Z. Ben-Shaul, Gail E. Kaiser, and George T. Heineman. An architecture for multiuser software development environments. Computing Systems, 6 (2):65–103, Spring 1993. CODEN CMSYE2. ISSN 0895-6340.

Badger:COMPSYS-9-1-47

[BSSW96] Lee Badger, Daniel F. Sterne, David L. Sherman, and Kenneth M. Walker. A domain and type enforcement UNIX prototype. Computing Systems, 9(1): 47–83, Winter 1996. CODEN CMSYE2. ISSN 0895-6340.

Burgess:COMPSYS-8-3-309

[Bur95] Mark Burgess. A site configuration engine. Computing Systems, 8(3):309–337, Summer 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Barton:COMPSYS-1-2-111

[BW88] J. M. Barton and J. C. Wagner. Enhanced resource sharing in UNIX. Computing Systems, 1 (2):111–133, Spring 1988. CO-DEN CMSYE2. ISSN 0895-6340.

Bhargava:COMPSYS-4-3-277

[BZM91] Bharat Bhargava, Yonguang Zhang, and Enrique Mafla. Evolution of a communication system for distributed transaction processing in Raid. Computing Systems, 4(3):277–313, Summer 1991. CODEN CMSYE2. ISSN 0895-6340.

Calabrese:COMPSYS-9-4-411

[Cal96a] Christopher J. Calabrese. Corrigendum: “A tool for building firewall-router configurations”. Computing Systems, 9(4):411, Fall 1996. CODEN CMSYE2. ISSN 0895-6340. See [Cal96b].

Calabrese:COMPSYS-9-3-239

[Cal96b] Christopher J. Calabrese. A tool for building firewall-router configurations. Computing Systems, 9(3):239–253, Summer 1996. CO-DEN CMSYE2. ISSN 0895-6340. See corrigendum [Cal96a].

Cargill:COMPSYS-4-1-69

[Car91] Tom A. Cargill. Controversy: The case against multiple inheritance in C++. Computing Systems, 4(1):69–82, Winter 1991. CODEN CMSYE2. ISSN 0895-6340. Cardelli:COMPSYS-8-1-27

[Car95]  Luca Cardelli.  A language

with distributed scope. Computing Systems, 8(1):27–59, Winter 1995. CODEN CMSYE2. ISSN 0895-6340.

Chen:COMPSYS-6-3-207

[CBE93] Jiansen Chen, Omran A. Buhkres, and Ahmed K. Elmagarmid. The implementation of cooperative mechanisms among system components in a heterogeneous multidatabase environment. Computing Systems, 6(3): 207–243, Summer 1993. CODEN CMSYE2. ISSN 0895-6340.

Casas:COMPSYS-8-2-171

[CCK^{+}95] Jeremy Casas, Dan L. Clark, Ravi Konuru, Steve W. Otto, Robert M. Prouty, and Jonathan Walpole. MPVM: A migration transparent version of PVM. Computing Systems, 8(2):171– 216, Spring 1995. CODEN CM-SYE2. ISSN 0895-6340. Comer:COMPSYS-3-4-487

[CDM90] Douglas Comer, Ralph E. Droms, and Thomas P. Murtagh. An experimental implementation of the tilde naming system. Computing Systems, 3(4):487–515, Fall 1990. CODEN CMSYE2. ISSN 0895-6340.

Campbell:COMPSYS-6-4-363

[CI93] Roy H. Campbell and Nayeem Islam. A technique for documenting the framework of an objectoriented system. Computing Systems, 6(4):363–389, Fall 1993. CODEN CMSYE2. ISSN 0895-6340. Campbell:COMPSYS-5-3-217

[CIM92] Roy H. Campbell, Nayeem Islam, and Peter Madany. Choices, frameworks and refinement. Computing Systems, 5(3):217– 257, Summer 1992. CODEN CM-SYE2. ISSN 0895-6340. Cheriton:COMPSYS-9-3-179

[CK96] David R. Cheriton and Robert A.

Kutter. Optimized memorybased messaging: Leveraging the memory system for high-performance communication. Computing Systems, 9(3): 179–215, Summer 1996. CODEN CMSYE2. ISSN 0895-6340. Cabrera:COMPSYS-4-4-405

[CL91] Luis-Felipe F. Cabrera and Darrell D. E. Long. Swift: Using distributed disk striping to provide high I/O data rates. Computing Systems, 4(4):405–436, Fall 1991. CODEN CMSYE2. ISSN 0895-6340. Cabrera:COMPSYS-5-3-199

[CLS92] Luis-Felipe Cabrera, Allen W.

Luniewski, and James W. Stamos. Fine-grained access control in a transactional object-oriented system. Computing Systems, 5 (3):199–216, Summer 1992. CO-DEN CMSYE2. ISSN 0895-6340.

Collyer:COMPSYS-9-2-119

[Col96] Geoff Collyer. Setting interrupt priorities in software via interrupt queueing. Computing Systems, 9 (2):119–130, Spring 1996. CO-DEN CMSYE2. ISSN 0895-6340.

Curran:COMPSYS-3-4-551

[CS90] Stephan Curran and Michael Stumm. A comparison of basic CPU scheduling algorithms for multiprocessor UNIX. Computing Systems, 3(4):551–579, Fall 1990. CODEN CMSYE2. ISSN 0895-6340.

Carson:COMPSYS-7-2-263

[CS94] Scott Carson and Sanjeev Setia. Optimal write batch size in log-structured file systems. Computing Systems, 7(2):263– 281, Spring 1994. CODEN CM-SYE2. ISSN 0895-6340.

Dasgupta:COMPSYS-4-3-243

[DAM^{+}91] Partha Dasgupta, R. Ananthanarayanan, Sathis Menon, Ajay Mohindra, and Raymond Chen. Distributed programming with objects and threads in the Clouds system. Computing Systems, 4 (3):243–275, Summer 1991. CO-DEN CMSYE2. ISSN 0895-6340.

Dasgupta:COMPSYS-3-1-11

[DCM^{+}90] P. Dasgupta, R. C. Chen, S. Menon, M. P. Pearson, R. Ananthanarayanan, U. Ramachandran, M. Ahamad, R. J. LeBlanc, W. F. Appelbe, J. M. Bernabéu-Aubán, P. W. Hutto, M. Y. A. Khalidi, and C. J.

Wilkenloh. The design and implementation of the Clouds distributed operating system. Computing Systems, 3(1):11–46, Winter 1990. CODEN CMSYE2. ISSN 0895-6340.

Dearle:COMPSYS-7-3-289

[DdBF^{+}94] Alan Dearle, Rex di Bona, James Farrow, Frans Kenskens, Anders Lindström, John Rosenberg, and Francis Vaughan. Grasshopper: An orthogonally persistent operating system. Computing Systems, 7(3):289–312, Summer 1994. CODEN CMSYE2. ISSN 0895-6340.

Deutsch:COMPSYS-5-4-375

[Deu92] Peter Deutsch. Guest editorial. Computing Systems, 5(4): 375–378, Fall 1992. CODEN CM-SYE2. ISSN 0895-6340.

Dewan:COMPSYS-6-2-61

[Dew93] Prasun Dewan. Guest editorial.

Computing Systems, 6(2):61–63, Spring 1993. CODEN CMSYE2. ISSN 0895-6340.

Davis:COMPSYS-9-1-29

[DGT96] Donald T. Davis, Daniel E. Geer, and Theodore Ts’o. Kerberos with clocks adrift: History, protocols, and implementation. Computing Systems, 9(1):29–46, Winter 1996. CODEN CMSYE2. ISSN 0895-6340.

Donner:COMPSYS-1-1-33

[DJ88] Marc D. Donner and David H.

Jameson. Language and operating system features for real-time programming. Computing Systems, 1(1):33–62, Winter 1988. CODEN CMSYE2. ISSN 0895-6340. Douglis:COMPSYS-8-4-381

[DK95] Fred Douglis and P. Krishnan. Adaptive disk spin-down policies for mobile computers. Computing Systems, 8(4):381–413, Fall 1995. CODEN CMSYE2. ISSN 0895-6340. Danzig:COMPSYS-5-4-433

[DLO92] Peter B. Danzig, Shih-Hao Li, and Katia Obraczka. Distributed indexing of autonomous Internet services. Computing Systems, 5 (4):433–459, Fall 1992. CODEN CMSYE2. ISSN 0895-6340. Douglis:COMPSYS-4-4-353

[DOKT91] Fred Douglis, John K. Ousterhout, M. Frans Kaashoek, and Andrew S. Tanenbaum. A comparison of two distributed systems: Amoeba and sprite. Computing Systems, 4(4):353– 384, Fall 1991. CODEN CM-SYE2. ISSN 0895-6340.

Duff:COMPSYS-2-2-155

[Duf89] Tom Duff. Experience with viruses on UNIX systems. Computing Systems, 2(2):155–171, Spring 1989. CODEN CMSYE2. ISSN 0895-6340. Dewan:COMPSYS-3-4-517

[DV90] Prasun Dewan and Eric Vasilik. An object model for conventional operating systems. Computing Systems, 3(4):517–549, Fall 1990. CODEN CMSYE2. ISSN 0895-6340.

Feng:COMPSYS-9-3-217

[FC96] Li-Chi Feng and Ruei-Chuan Chang. Achieving predictable response time with an intelligent file system updater. Computing Systems, 9(3):217–237, Summer 1996. CODEN CMSYE2. ISSN 0895-6340. Feldman:COMPSYS-3-2-359

[FG90] Stuart Feldman and W. Morven Gentleman. Portability— A no longer solved problem. Computing Systems, 3(2):359– 380, Spring 1990. CODEN CM-SYE2. ISSN 0895-6340. Golding:COMPSYS-5-4-379

[Gol92] Richard A. Golding. A weakconsistency architecture for distributed information services. Computing Systems, 5(4):379– 405, Fall 1992. CODEN CM-SYE2. ISSN 0895-6340.

Grass:COMPSYS-5-1-5

[Gra92] Judith E. Grass. Object-oriented design archaeology with CIA++. Computing Systems, 5(1):5–67, Winter 1992. CODEN CMSYE2. ISSN 0895-6340. Griswold:COMPSYS-2-4-339

[Gri89] Ralph E. Griswold. Data structures in the icon programming language. Computing Systems, 2 (4):339–365, Fall 1989. CODEN CMSYE2. ISSN 0895-6340. Hawley:COMPSYS-3-2-289

[Haw90] Michael Hawley. The personal orchestra, or audio data compression by 10,000:1. Computing Systems, 3(2):289–329, Spring 1990.

CODEN CMSYE2. ISSN 0895-6340.

Herrin:COMPSYS-4-2-127

[HF91] Erin H. Herrin II and Raphael Finkel. An ASCII database for fast queries of relatively stable data. Computing Systems, 4(2): 127–155, Spring 1991. CODEN CMSYE2. ISSN 0895-6340.

Herrin:COMPSYS-9-2-93

[HF96] Eric H. Herrin, II and Raphael A. Finkel. Schema and tuple trees: An intuitive structure for representing relational data. Computing Systems, 9(2):93–118, Spring 1996. CODEN CMSYE2. ISSN 0895-6340.

Huston:COMPSYS-8-4-365

[HH95] L. B. Huston and P. Honeyman. Partially connected operation. Computing Systems, 8(4): 365–379, Fall 1995. CODEN CM-SYE2. ISSN 0895-6340.

Honeyman:COMPSYS-7-2-173

[Hon94] Peter Honeyman. Guest editorial. Computing Systems, 7(2): 173, Spring 1994. CODEN CM-SYE2. ISSN 0895-6340.

Heydon:COMPSYS-7-1-91

[HT94] Allen Heydon and J. D. Tygar. Specifying and checking UNIX security constraints. Computing Systems, 7(1):91–112, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Ingham:COMPSYS-7-3-313

[IP94] David B. Ingham and Graham D. Parrington. Delayline:

A wide-area network emulation tool. Computing Systems, 7(3): 313–332, Summer 1994. CODEN CMSYE2. ISSN 0895-6340.

Johnson:COMPSYS-1-2-159

[Joh88] Stephen C. Johnson. YACC meets C++. Computing Systems, 1(2):159–167, Spring 1988. CO-DEN CMSYE2. ISSN 0895-6340.

Johnson:COMPSYS-6-1-29

[Joh93] Steve Johnson. United we fall, or killing the goose that laid the golden egg. Computing Systems, 6(1):29–34, Winter 1993. CO-DEN CMSYE2. ISSN 0895-6340.

Kordale:COMPSYS-9-4-377

[KAD96] R. Kordale, M. Ahamad, and M. Devarakonda. Object caching in a CORBA compliant system. Computing Systems, 9(4): 377–404, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Kotz:COMPSYS-9-3-155

[KC96] David Kotz and Preston Crow.

The expected lifetime of singleaddress-space operating systems. Computing Systems, 9(3):155– 178, Summer 1996. CODEN CM-SYE2. ISSN 0895-6340.

Krajewski:COMPSYS-7-1-75

[KCCT94] Marjan Krajewski, Jr., John C.

Chipchak, David A. Chodorow, and Jonathan T. Trostle. Applicability of smart cards to network user authentication. Computing Systems, 7(1):75–89, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Knister:COMPSYS-6-2-135

[KP93] Michael Knister and Atul Prakash. Issues in the design of a toolkit for supporting multiple group editors. Computing Systems, 6(2):135–166, Spring 1993. CODEN CMSYE2. ISSN 0895-6340.

Karsenty:COMPSYS-6-2-167

[KTBL93] Alain Karsenty, Christophe Tronche, and Michel Beaudouin-Lafon. GroupDesign: Shared editing in a heterogeneous environment. Computing Systems, 6 (2):167–195, Spring 1993. CO-DEN CMSYE2. ISSN 0895-6340.

Kernighan:COMPSYS-2-2-103

[KV89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. Computing Systems, 2 (2):103–132, Spring 1989. CO-DEN CMSYE2. ISSN 0895-6340.

URL http://www.usenix.org/
publications/compsystems/1989/
spr_kernighan.pdf.

Langston:COMPSYS-3-2-193

[Lan90] Peter S. Langston. Little languages for music. Computing Systems, 3(2):193–288, Spring 1990. CODEN CMSYE2. ISSN 0895-6340.

LaPadula:COMPSYS-7-1-113

[LaP94] Leonard J. LaPadula. A ruleset approach to formal modeling of a trusted computer system. Computing Systems, 7(1): 113–167, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Laumann:COMPSYS-7-4-419

[LB94] Oliver Laumann and Carsten Bormann. Elk: The Extension Language Kit. Computing Systems, 7(4):419–449, Fall 1994. CODEN CMSYE2. ISSN 0895-6340.

Lesk:COMPSYS-1-2-189

[Les88a] Michael Lesk. Can Unix survive secret source code? Computing Systems, 1(2):189–199, Spring 1988. CODEN CMSYE2. ISSN 0895-6340.

Lesk:COMPSYS-1-3-207

[Les88b] Michael Lesk. GRAB—inverted indexes with low storage overhead. Computing Systems, 1(3): 207–220, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Libes:COMPSYS-4-2-99

[Lib91] Don Libes. expect: Scripts for controlling interactive processes. Computing Systems, 4(2):99–125, Spring 1991. CODEN CMSYE2. ISSN 0895-6340.

Long:COMPSYS-7-3-333

[LM94] Darrell D. E. Long and Bruce R.

Montague. Swift/RAID: A distributed RAID system. Computing Systems, 7(3):333–359, Summer 1994. CODEN CMSYE2. ISSN 0895-6340.

McIlroy:COMPSYS-6-1-5

[MBM93] Peter M. McIlroy, Keith Bostic, and M. Douglas McIlroy. Engineering radix sort. Computing Systems, 6(1):5–27, Winter 1993. CODEN CMSYE2. ISSN 0895-

McIlroy:COMPSYS-2-2-173

[McI89] M. Douglas McIlroy. Virology 101 (UNIX system virus). Computing Systems, 2(2):173–181, Spring 1989. CODEN CMSYE2. ISSN 0895-6340.

McKusick:COMPSYS-8-1-3

[McK95] Marshall Kirk McKusick. The virtual filesystem interface in 4.4BSD. Computing Systems, 8 (1):3–25, Winter 1995. CODEN CMSYE2. ISSN 0895-6340.

McKenney:COMPSYS-5-2-141

[MD92] Paul E. McKenney and Ken F. Dove. Efficient demultiplexing of incoming TCP packets. Computing Systems, 5(2):141– 157, Spring 1992. CODEN CM-SYE2. ISSN 0895-6340.

Mock:COMPSYS-5-3-259

[MKC92] Michael Mock, Reinhold Kroeger, and Vinny Cahill. Implementing atomic objects with the RelaX transaction facility. Computing Systems, 5(3):259–304, Summer 1992. CODEN CMSYE2. ISSN 0895-6340.

Moffat:COMPSYS-5-4-494

[Mof92a] Alistair Moffat. Corrigendum. Computing Systems, 5(4):494, Fall 1992. CODEN CMSYE2. ISSN 0895-6340. See [Mof92b].

Moffat:COMPSYS-5-2-125

[Mof92b] Alistair Moffat. Economical inversion of large text files. Computing Systems, 5(2):125–139, Spring 1992. CODEN CMSYE2.

ISSN 0895-6340. See corrigendum in [Mof92a].

Mogul:COMPSYS-7-2-201

[Mog94] Jeffrey C. Mogul. Recovery in spritely NFS. Computing Systems, 7(2):201–262, Spring 1994. CODEN CMSYE2. ISSN 0895-6340.

Morris:COMPSYS-1-4-405

[Mor88] Robert A. Morris. An unorthodox approach to undergraduate software engineering instruction. Computing Systems, 1(4): 405–419, Fall 1988. CODEN CM-SYE2. ISSN 0895-6340.

Massalin:COMPSYS-3-1-139

[MP90] Henry Massalin and Calton Pu.

Fine-grain adaptive scheduling using feedback. Computing Systems, 3(1):139–173, Winter 1990. CODEN CMSYE2. ISSN 0895-6340.

Nemeth:COMPSYS-1-1-3

[Nem88] Alan G. Nemeth. Welcome. Computing Systems, 1(1):3, Winter 1988. CODEN CMSYE2. ISSN 0895-6340.

Neuman:COMPSYS-5-4-407

[Neu92] B. Clifford Neuman. The Prospero file system: A global file system based on the virtual system model. Computing Systems, 5(4): 407–432, Fall 1992. CODEN CM-SYE2. ISSN 0895-6340.

Nilsen:COMPSYS-7-4-467

[Nil94] Kevin Nilsen. Reliable realtime garbage collection of C++.

Computing Systems, 7(4):467– 504, Fall 1994. CODEN CM-SYE2. ISSN 0895-6340.

Noble:COMPSYS-8-4-345

[NPS95] Brian D. Noble, Morgan Price, and M. Satyanarayanan. A programming interface for applicationaware adaptation in mobile computing. Computing Systems, 8(4): 345–363, Fall 1995. CODEN CM-SYE2. ISSN 0895-6340.

Odell:COMPSYS-1-1-5

[O’D88a] Michael O’Dell. Greetings. Computing Systems, 1(1):5–9, Winter 1988. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-1-2-107

[O’D88b] Michael D. O’Dell. Greetings. Computing Systems, 1(2):107– 109, Spring 1988. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-1-4-301

[O’D88c] Michael D. O’Dell. Greetings. Computing Systems, 1(4):301– 304, Fall 1988. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-1-3-205

[O’D88d] Michael D. O’Dell. Greetings and felicitations. Computing Systems, 1(3):205–206, Summer 1988. CO-DEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-2-1-3

[O’D89a] Michael D. O’Dell. Greetings. Computing Systems, 2(1): 3–5, Winter 1989. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-2-2-99

[O’D89b] Michael D. O’Dell. Greetings.

Computing Systems, 2(2):99–101, Spring 1989. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-2-3-189

[O’D89c] Michael D. O’Dell. Greetings.

Computing Systems, 2(3):189– 190, Summer 1989. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-2-4-285

[O’D89d] Michael D. O’Dell. Greetings. Computing Systems, 2(4): 285, Fall 1989. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-3-2-191

[O’D90a] Michael D. O’Dell. Greetings.

Computing Systems, 3(2):191– 192, Spring 1990. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-3-3-385

[O’D90b] Michael D. O’Dell. Greetings.

Computing Systems, 3(3):385– 386, Summer 1990. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-3-4-483

[O’D90c] Michael D. O’Dell. Greetings.

Computing Systems, 3(4):483– 484, Fall 1990. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-4-1-3

[O’D91a] Michael D. O’Dell. Greetings. Computing Systems, 4(1): 3–4, Winter 1991. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-4-2-97

[O’D91b] Michael D. O’Dell. Greetings. Computing Systems, 4(2):97–98, Spring 1991. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-4-4-351

[O’D91c] Michael D. O’Dell. Greetings! Computing Systems, 4(4):351– 352, Fall 1991. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-5-4-373

[O’D92a] Michael O’Dell. Greetings. Computing Systems, 5(4):373–374, Fall 1992. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-5-1-3

[O’D92b] Michael D. O’Dell. Greetings! Computing Systems, 5(1):3, Winter 1992. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-5-2-123

[O’D92c] Michael D. O’Dell. Greetings. Computing Systems, 5(2):123– 124, Spring 1992. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-5-3-197

[O’D92d] Michael D. O’Dell. Greetings. Computing Systems, 5(3): 197, Summer 1992. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-6-1-3

[O’D93a] Michael D. O’Dell. Greetings. Computing Systems, 6(1): 3–4, Winter 1993. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-6-4-361

[O’D93b] Michael D. O’Dell. Greetings.

Computing Systems, 6(4):361– 362, Fall 1993. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-7-3-285

[O’D94a] Michael D. O’Dell. Greetings.

Computing Systems, 7(3):285– 286, Summer 1994. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-7-4-415

[O’D94b] Michael D. O’Dell. Greetings, dear friends, one last time. Computing Systems, 7(4):415– 417, Fall 1994. CODEN CM-SYE2. ISSN 0895-6340.

ODell:COMPSYS-3-1-3

[OS89] Michael D. O’Dell and Peter H.

Salus. Greetings. Computing Systems, 3(1):3, Winter 1989. CO-DEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-4-3-177

[OS91] Michael D. O’Dell and Peter H.

Salus. Greetings. Computing Systems, 4(3):177, Summer 1991. CODEN CMSYE2. ISSN 0895-6340.

ODell:COMPSYS-7-1-iii

[OS94] Michael D. O’Dell and Peter H.

Salus. Special greetings. Computing Systems, 7(1):iii, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Oyang:COMPSYS-7-3-393

[OW94] Yen-Jen Oyang and Le-Chun Wu. Optimal design of megabyte second-level caches for minimizing bus traffic in sharedmemory shared-bus multiprocessors. Computing Systems, 7(3): 393–408, Summer 1994. CODEN CMSYE2. ISSN 0895-6340.

Pucci:COMPSYS-3-1-47

[PA90] Marc F. Pucci and J. L. Alberi. Using hints in DUNE remote procedure calls. Computing Systems, 3(1):47–68, Winter 1990. CO-DEN CMSYE2. ISSN 0895-6340.

Parrington:COMPSYS-8-2-135

[Par95]  Graham D. Parrington.  A

stub generation system for C++. Computing Systems, 8(2):135– 169, Spring 1995. CODEN CM-SYE2. ISSN 0895-6340.

Pyarali:COMPSYS-9-4-331

[PHS96] Irfan Pyarali, Timothy H. Harrison, and Douglas C. Schmidt. Design and performance of an object-oriented framework for high-speed electronic medical imaging. Computing Systems, 9 (4):331–375, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Pike:COMPSYS-1-3-279

[Pik88] Rob Pike. Controversy: Window systems should be transparent. Computing Systems, 1(3): 279–296, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Pike:COMPSYS-2-2-133

[Pik89] Rob Pike. A concurrent window system. Computing Systems, 2 (2):133–153, Spring 1989. CO-DEN CMSYE2. ISSN 0895-6340.

Patel:COMPSYS-6-2-105

[PK93] Dorab Patel and Scott D. Kalter.

A UNIX toolkit for distributed synchronous collaborative applications. Computing Systems, 6 (2):105–133, Spring 1993. CO-DEN CMSYE2. ISSN 0895-6340.

Pu:COMPSYS-1-1-11

[PMI88] Calton Pu, Henry Massalin, and John Ioannidis. The synthesis kernel. Computing Systems, 1 (1):11–32, Winter 1988. CODEN CMSYE2. ISSN 0895-6340.

Pike:COMPSYS-8-3-221

[PPD^{+}95] Rob Pike, Dave Presotto, Sean Dorward, Bob Flandrena, Ken Thompson, Howard Trickey, and Phil Winterbottom. Plan 9 from Bell Labs. Computing Systems, 8 (3):221–254, Summer 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Presotto:COMPSYS-8-1-1

[Pre95] David L. Presotto. Editorial.

Computing Systems, 8(1):1, Winter 1995. CODEN CMSYE2. ISSN 0895-6340.

Parrington:COMPSYS-8-3-255

[PSWL95] Graham D. Parrington, Santosh K. Shrivastava, Stuart M. Wheater, and Mark C. Little. The design and implementation of Arjuna. Computing Systems, 8 (3):255–308, Summer 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Pucci:COMPSYS-4-3-217

[Puc91] Marc F. Pucci. Configurable data manipulation in an attached multiprocessor. Computing Systems, 4(3):217–242, Summer 1991. CO-DEN CMSYE2. ISSN 0895-6340.

Rozier:COMPSYS-1-4-305

[RAA^{+}88] M. Rozier, V. Abrossimov, F. Armand, I. Boule, M. Gien, M. Guillemont, F. Herrmann, C. Kaiser, S. Langlois, P. Lé oñard, and W. Neuhauser. CHO-RUS distributed operating system. Computing Systems, 1(4): 305–370, Fall 1988. CODEN CM-SYE2. ISSN 0895-6340.

Rees:COMPSYS-8-4-343

[Ree95] Jim Rees. Guest editor. Computing Systems, 8(4):343–344, Fall 1995. CODEN CMSYE2. ISSN 0895-6340.

Rosenberg:COMPSYS-5-3-305

[Ros92] John Rosenberg. Architectural and operating system support for orthogonal persistence. Computing Systems, 5(3):305–335, Summer 1992. CODEN CMSYE2. ISSN 0895-6340.

Ruane:COMPSYS-3-3-387

[Rua90] Lawrence M. Ruane. Process synchronization in the UTS kernel. Computing Systems, 3(3): 387–421, Summer 1990. CODEN CMSYE2. ISSN 0895-6340.

Rubin:COMPSYS-9-1-15

[Rub96a] Aviel D. Rubin. Independent one-time passwords. Computing Systems, 9(1):15–27, Winter 1996. CODEN CMSYE2. ISSN 0895-6340. See also [Ada96, Rub96b].

Rubin:COMPSYS-9-3-257

[Rub96b] Aviel D. Rubin. Response: Independent one-time passwords. Computing Systems, 9(3):257, Summer 1996. CODEN CM-SYE2. ISSN 0895-6340. See [Rub96a, Ada96].

Riggs:COMPSYS-9-4-291

[RWWB96] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. Computing Systems, 9(4): 291–312, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Sakkinen:COMPSYS-5-3-361

[Sak92a] Markku Sakkinen. Corrigendum to A Critique of the Inheritance Principles of C++. Computing Systems, 5(3):361–363, Summer 1992. CODEN CMSYE2. ISSN 0895-6340.

Sakkinen:COMPSYS-5-1-69

[Sak92b] Markku Sakkinen. A critique of the inheritance principles of C++. Computing Systems, 5(1): 69–110, Winter 1992. CODEN CMSYE2. ISSN 0895-6340.

Salus:COMPSYS-3-4-485

[Sal90] Peter H. Salus. Tom strong (1944–1990). Computing Systems, 3(4):485, Fall 1990. CO-DEN CMSYE2. ISSN 0895-6340.

Salus:COMPSYS-7-3-287

[Sal94]  Peter H. Salus.  Announce-

ments. Computing Systems, 7(3): 287, Summer 1994. CODEN CM-SYE2. ISSN 0895-6340.

Salus:COMPSYS-8-3-iii

[Sal95a] Peter H. Salus. A note from the managing editor. Computing Systems, 8(3):iii, Summer 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Salus:COMPSYS-8-4-341

[Sal95b] Peter H. Salus. A note from the managing editor. Computing Systems, 8(4):341, Fall 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Salus:COMPSYS-9-4-405

[Sal96] Peter H. Salus. Exeunt omnes. Computing Systems, 9(4):405, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Schmidt:COMPSYS-9-4-261

[Sch96] Douglas C. Schmidt. Guest editorial. Computing Systems, 9(4): 261–263, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Schwartz:COMPSYS-5-4-461

[SEKN92] Michael F. Schwartz, Alan Emtage, Brewster Kahle, and B. Clifford Neuman. A comparison of Internet resource discovery approaches. Computing Systems, 5(4):461–493, Fall 1992. CODEN CMSYE2. ISSN 0895-6340.

Shapiro:COMPSYS-2-4-287

[SGH^{+}89] Marc Shapiro, Yvon Gourbant, Sabine Habert, Laurence Mosseri, Michel Ruffin, and Celine Valot. SOS: An object-oriented operating system —assessment and perspectives. Computing Systems, 2(4):287– 337, Fall 1989. CODEN CM-SYE2. ISSN 0895-6340.

Scott:COMPSYS-3-1-101

[SLM^{+}90] Michael L. Scott, Thomas J.

LeBlanc, Brian D. Marsh, Timothy G. Becker, Cezary Dubnicki, Evangelos P. Markatos, and Neil G. Smithline. Implementation issues for the Psyche multiprocessor operating system. Computing Systems, 3(1): 101–138, Winter 1990. CODEN CMSYE2. ISSN 0895-6340.

Smith:COMPSYS-1-3-255

[SM88] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. Computing Systems, 1(3):255–278, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Smith:COMPSYS-4-4-385

[Smi91] Jonathan M. Smith. The software design laboratory. Computing Systems, 4(4):385–404, Fall 1991. CODEN CMSYE2. ISSN 0895-6340.

Sosic:COMPSYS-8-2-107

[Sos95] Rok Sosič. The Dynascope Directing Server: Design and implementation. Computing Systems, 8(2):107–134, Spring 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Spafford:COMPSYS-3-1-5

[Spa89] Eugene H. Spafford. Guest editorial. Computing Systems, 3(1): 5–9, Winter 1989. CODEN CM-SYE2. ISSN 0895-6340.

Spafford:COMPSYS-4-3-179

[Spa91] Eugene Spafford. Guest editorial. Computing Systems, 4(3): 179–183, Summer 1991. CODEN CMSYE2. ISSN 0895-6340.

Srinidhi:COMPSYS-6-3-285

[Sri93] Hassan N. Srinidhi. Managing data redundancy in interoperable heterogeneous environments. Computing Systems, 6(3): 285–317, Summer 1993. CODEN CMSYE2. ISSN 0895-6340.

Sousa:COMPSYS-6-4-391

[SSZ^{+}93] Pedro Sousa, Manuel Sequeira, André Zúquete, Paulo Ferreira, Cristina Lopes, José Pereira, Paulo Guedes, and José Alves Marques. Distribution and persistence in the IK platform: Overview and evaluation. Computing Systems, 6(4):391–424, Fall 1993. CODEN CMSYE2. ISSN 0895-6340.

Stevens:COMPSYS-2-3-251

[Ste89] W. Richard Stevens. Heuristics for disk drive positioning in 4.3 BSD. Computing Systems, 2(3): 251–274, Summer 1989. CODEN CMSYE2. ISSN 0895-6340.

Stroustrup:COMPSYS-1-4-371

[Str88] Bjarne Stroustrup. Type-safe linkage for C++. Computing Systems, 1(4):371–403, Fall 1988. CODEN CMSYE2. ISSN 0895-6340.

Stroustrup:COMPSYS-2-3-191

[Str89a] Bjarne Stroustrup. The evolution of C++: 1985 to 1989. Computing Systems, 2(3):191–250, Summer 1989. CODEN CMSYE2. ISSN 0895-6340.

Stroustrup:COMPSYS-2-4-367

[Str89b] Bjarne Stroustrup. Multiple inheritance for C++. Computing Systems, 2(4):367–395, Fall 1989. CODEN CMSYE2. ISSN 0895-6340.

Stroustrup:COMPSYS-2-1-55

[Str89c] Bjarne Stroustrup. Parametrized types for C++. Computing Systems, 2(1):55–85, Winter 1989. CODEN CMSYE2. ISSN 0895-6340.

Spezzano:COMPSYS-3-3-423

[STV90] G. Spezzano, D. Talia, and M. Vanneschi. A concurrent programming support for distributed systems. Computing Systems, 3 (3):423–447, Summer 1990. CO-DEN CMSYE2. ISSN 0895-6340.

Thompson:COMPSYS-3-2-331

[Tho90] Tim Thompson. Keynote—A language and extensible graphics editor for music. Computing Systems, 3(2):331–357, Spring 1990. CODEN CMSYE2. ISSN 0895-6340.

Urban:COMPSYS-6-3-203

[UB93] Susan Urban and Elisa Bertino.

Guest editorial. Computing Systems, 6(3):203–206, Summer 1993. CODEN CMSYE2. ISSN 0895-6340.

Viles:COMPSYS-8-1-61

[VF95] Charles L. Viles and James C.

French. Availability and latency of World Wide Web information servers. Computing Systems, 8 (1):61–91, Winter 1995. CODEN CMSYE2. ISSN 0895-6340.

Vaughan:COMPSYS-5-3-337

[VLD^{+}92] Francis Vaughan, Tracy Lo Basso, Alan Dearle, Chris Marlin, and Chris Barter. Casper: a cached architecture supporting persistence. Computing Systems, 5(3):337–359, Summer 1992. CO-DEN CMSYE2. ISSN 0895-6340.

Wagner:COMPSYS-3-3-449

[Wag90] Bernhard Wagner. Distributed spooling in a heterogeneous environment. Computing Systems, 3 (3):449–477, Summer 1990. CO-DEN CMSYE2. ISSN 0895-6340.

Waldo:COMPSYS-4-2-157

[Wal91] Jim Waldo. Controversy: The case for multiple inheritance in C++. Computing Systems, 4(2): 157–171, Spring 1991. CODEN CMSYE2. ISSN 0895-6340.

Ware:COMPSYS-7-1-1

[War94] Willis H. Ware. Policy considerations for data networks. Computing Systems, 7(1):1–44, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Whang:COMPSYS-6-3-319

[WCN93] Whan-Kyu Whang, Sharma Chakravarthy, and Shamkant B. Navathe. Relational schema integration: Dealing with interrelation correspondences and querying over component relations. Computing Systems, 6(3):

319–352, Summer 1993. CODEN CMSYE2. ISSN 0895-6340.

Welch:COMPSYS-4-3-315

[Wel91] Brent B. Welch. Measured performance of caching in the Sprite network file system. Computing Systems, 4(3):315–342, Summer 1991. CODEN CMSYE2. ISSN 0895-6340.

Welch:COMPSYS-7-2-175

[Wel94] Brent Welch. A comparison of three distributed file system architectures: Vnode, Sprite, and Plan 9. Computing Systems, 7(2): 175–199, Spring 1994. CODEN CMSYE2. ISSN 0895-6340.

Wilhelms:COMPSYS-1-1-63

[Wil88a] Jane Wilhelms. Dynamics for computer graphics: A tutorial. Computing Systems, 1(1):63–94, Winter 1988. CODEN CM-SYE2. ISSN 0895-6340. See also [Wil88b].

Wilhelms:COMPSYS-1-2-200

[Wil88b] Jane Wilhelms. Erratum in “Dynamics for Computer Graphics”. Computing Systems, 1(2): 200, Spring 1988. CODEN CM-SYE2. ISSN 0895-6340. See [Wil88a].

Winckler:COMPSYS-7-3-361

[Win94]  Andreas Winckler.  A dis-

tributed look-ahead workload assignment algorithm for interdependent tasks. Computing Systems, 7(3):361–391, Summer 1994. CODEN CMSYE2. ISSN 0895-6340.

Winkler:COMPSYS-9-1-3

[Win96] Ira S. Winkler. The non-technical threat to computing systems. Computing Systems, 9(1):3–14, Winter 1996. CODEN CMSYE2. ISSN 0895-6340.

Wollrath:COMPSYS-9-4-265

[WRW96] Ann Wollrath, Roger Riggs, and Jim Waldo. A distributed object model for the Java system. Computing Systems, 9(4): 265–290, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Yahalom:COMPSYS-7-4-451

[Yah94] Raphael Yahalom. Secure timeliness: On the cost of nonsynchronized clocks. Computing Systems, 7(4):451–465, Fall 1994. CODEN CMSYE2. ISSN 0895-6340.

Yahalom:COMPSYS-7-1-45

[YKB94] Raphael Yahalom, Birgit Klein, and Thomas Beth. Trust-based navigation in distribution systems. Computing Systems, 7(1): 45–73, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.